OpenStudy (anonymous):

For his phone service, Kevin pays a monthly fee of $12 , and he pays an additional $0.06 per minute of use. The least he has been charged in a month is $99.06. What are the possible numbers of minutes he has used his phone in a month? use m for the number of minutes, and solve your inequality for m

OpenStudy (anonymous):

You're told here that the monthly bill is the monthly fee + $0.06 times the number of minutes. If m is the number of minutes. What algebraic expression represents the monthly bill? :-)

OpenStudy (anonymous):

the 12 dollars represents the monthly bill, so 12+0.06

OpenStudy (anonymous):

Close; you're missing a very important part of that statement. It's 0.06 per minute. So how much are you charged for m minutes, given that it costs 0.06 per minute?

OpenStudy (anonymous):

erm, 12+0.06=99.06?

OpenStudy (anonymous):

No, it's \(12+0.06\color{red}m \geq 99.06\) (why is it greater than or equal to?)

OpenStudy (anonymous):

because it's the number of minutes he's been charged in the month? o.O

OpenStudy (anonymous):

It's because that was the LEAST amount he's ever been charged (this statement allows for the scenarios when his bill was more than 99.06). Does this clarify things?

OpenStudy (anonymous):

It says to write the answer in simplist form.. and yes it does clarify things :)

OpenStudy (anonymous):

Alright, now that we know \(12+0.06m\geq 99.06\), what do you get when you solve for \(m\)? :-)

OpenStudy (anonymous):

(Note that your answer will also be an inequality)

OpenStudy (anonymous):

well I subtracted 0.06 from both sides, am i doing okay so far?

OpenStudy (anonymous):

No, 0.06 is the coefficient of m; what you first want to do is subtract 12 from both sides. :-)

OpenStudy (anonymous):

is the answer 1451?

OpenStudy (anonymous):

That's part of it. What inequality did you end up with in terms of m?

OpenStudy (anonymous):

(i.e. did you get \(m\geq 1451\) or \(m\leq 1451\)?)

OpenStudy (anonymous):

m<1451

OpenStudy (anonymous):

Unfortunately no. :-/ You should have \(12+0.06m\geq 99.06 \implies m\geq 1451\).

OpenStudy (anonymous):

The interpretation of this answer is that he's billed for at least 1451 minutes every month. Does this make sense? :-)

OpenStudy (anonymous):

Yes it makes alot of sense :)I appreciate it